# Differentially Private Matchings

### v2

@@ 1.1.1 Solution Types
One natural approach is to allow our algorithm to output “edges” that are not true edges; in other words, our matching algorithm returns a matching where only some of the edges actually exist. This corresponds to changing the input graph $G=(V,E)$ to a weighted complete graph, where every $\{u,v\}\not\in E$ is assigned weight 00 in the complete graph and actual edges have weight $1$, and asking for a max-weight matching in this graph. We can even allow our algorithm to be bicriteria and return a $b$-matching [^1] Recall that in a $b$-matching every node has degree at most $b$, thus a traditional matching is a $1$-matching. rather than a true matching. In this case, we would want our returned $b$-matching to contain an (approximate) maximum matching, while minimizing the magnitude of $b$.
As our first result, we show a strong lower bound in this model, even if we allow ourselves to output a $b$-matching and only compare its weight to the maximum matching. More carefully, we define a $(\gamma,\beta)$-approximation to be an algorithm that outputs a $b$-matching whose total weight is at least $(\text{OPT}/\gamma)-\beta$, where OPT is the size of the maximum matching in $G$.
@@
$\gamma(e^{2\varepsilon}b+\delta n)+\beta\geq n/2.$
The proof of Theorem 1.1 is presented in Section 3. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then a few meaningful regimes of $b,\gamma,\beta$ include the following:
@@
  • If $\gamma\leq\frac{n}{cb}$ for some reasonably large constant $c$, then $\beta\geq\Omega(n)$. Thus we have linear additive loss even if we allow multiplicative loss of almost $n/b$.
We can actually strengthen this into a lower bound even on fractional matchings. A fractional $b$-matching is an assignment of values in $[0,1]$ to edges (i.e., a vector $X\in[0,1]^{\binom{V}{2}}$) so that $\sum_{v\in V}X_{\{u,v\}}\leq b$. As the proof of the following theorem proceeds similarly to to lower bound for explicit integral solutions, we defer it to Appendix B.
Theorem 1.2 (Lower Bound on Explicit Fractional Solutions).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a fractional $b$-matching $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even just in expectation) of the maximum matching of a $0/1$-weighted (Definition 2.7) input graph $G=(V,E)$, then
$\gamma(e^{2\varepsilon}b+\delta n)+\beta\geq n/2.$
Implicit Solutions.
@@
Theorem 1.1 suggests that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). That is, the information released to the billboard is $\varepsilon$-differentially private and each node computes its own part of the solution by combining the public information on the billboard with its own private adjacency list. Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 2.4; such a solution has been used previously in the context of vertex cover, set cover, and maximum matching in bipartite graphs [GuptaLMRT10, hsu2014private]. It is important to note that the billboard model is not a privacy model, but a solution release model.
1.1.2 Edge Privacy (LEDP)
@@
Our results for edge privacy are within the novel setting that is a combination of the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
@@
Although both [hsu2014private] and our results are in the billboard model for solution release, their algorithm hinges on the “binary mechanism” [CSS11], a mechanism for releasing a stream of $n$ private counting queries with only $O(\log n)$ error, rather than the $O(\operatorname{poly}(n))$ error attained through naive composition. This mechanism releases noisy partial sums such that each query can be answered by combining $O(\log n)$ partial sums. However, this requires a centralized curator who can compute the true partial sums before adding noise, which seems incompatible with the local privacy model.
In comparison, we can guarantee LEDP rather than centralized differential privacy, and hence our solutions are in a much stronger privacy model (although only for edge-privacy rather than for node-privacy). We also later provide solutions for node-DP in the centralized setting.
As with our lower bound in Theorem 1.1, we will allow our algorithms to be bicriteria and output matchings that have up to $b$-degree. Interestingly, we present three main results for matchings depending on the size of $b$.
Our solutions are additionally guaranteed to (approximate) maximum matchings, with high probability. Given a large enough gap between $b$ and $b^{\prime}$, we also show that our solutions are guaranteed to (approximate) maximum $b^{\prime}$-matchings. When $b$ is sufficiently large, our returned matching will be a $(2+\eta)$-approximation of the maximum $b$-matching, and would no longer need to be a bicriteria approximation.
Our first result is the following, which we show in Section 4.
Theorem 1.3 (LEDP $2$-Approximate Maximum Matching).
We further extend our result to $b^{\prime}$-matchings where given a sufficient gap between $b$ and $b^{\prime}$, we can obtain an $\varepsilon$-LEDP algorithm that outputs a matching with degree at most $b$ that is a $2$-approximate maximum $b^{\prime}$-matching. We present these results in Section 10.
@@
For $\varepsilon\in(0,1)$ and $\eta\in(0,1)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+O\left(\frac{\log(n)}{\eta^{2}\varepsilon}\right)$ that has the size of a $2$-approximate maximum $b^{\prime}$-matching.
Finally, when $b$ is sufficiently large, we no longer require a bicriteria approximation; instead, we are guaranteed to return an approximate maximum $b$-matching, with high probability. We present these results in detail in Section 10.
Theorem 1.5 (LEDP $2$-Approximate Maximum $b$-Matching).
@@
For $\varepsilon\in(0,1)$ and $\eta\in(0,1/2)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching, in the billboard model, for $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$ that has the size of a $(2+\eta)$-approximate maximum $b$-matching.
One downside of the algorithm used to prove Theorem 1.3 is its running time: it takes $\Omega(n)$ rounds. This is not ideal in the LEDP setting, where we have to wait for every node to communicate with the curator in each round before we can move onto the next, causing each round of the algorithm to be quite slow if there are straggler nodes. So we next show (in Section 5) an improved version that runs in only $O(\log n)$ rounds, based on combining the ideas of Theorem 1.3 with ideas from parallel and distributed algorithms. The price we pay for this efficiency is some extra loss in the $b$ parameter.
Theorem 1.6 (Efficient 2-Approximate Maximum Matching).
@@
For $\varepsilon\in(0,1)$, Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) matching with degree at most $b$ in the billboard model for $b=O(\log^{2}(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
@@
We complement our upper bounds with a lower bound which shows that even for the particular type of implicit solutions used in Theorem 1.3, our reliance on $b$ is necessary: if $b$ is sublogarithmic, then even an implicit solution cannot be a good approximation to the maximum matching in the graph. The following theorem is an informal version of this; see Section 11.1 and Theorem 11.1 for the precise statement.
@@
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 1.3 which is the size of a $\gamma$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\gamma$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\gamma$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\gamma$, then we say that $F$ is a $\gamma$-approximate $b$-matching. with probability at least $1-1/\mathrm{poly}(n)$. Then, when decoded from the billboard, with probability at least $1/2$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, regardless of the value of $\gamma>0$.
@@
1.1.3 Node Privacy
@@
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. It is a significantly more challenging setting than edge privacy. We show in Section 6 that the ideas we developed for edge-DP in general graphs can be combined with a “matching sparsifier” [solomon2021local] to give the first implicit node-DP algorithms for matching in general graphs. Specifically, we show the following upper bound results in node-DP within Section 6, where $\alpha$ is the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of the input graph.
Theorem 1.8 (Node-DP Approximate Maximum Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}\right)$, and (ii) the implicit solution has the size of the $(2+\eta)$-approximate maximum matching.
Our node-DP algorithms rely on a stable implementation of a bounded arboricity sparsifier that sparsifies the graph (in a manner that doesn’t increase edge edit distance between neighboring graphs) such that each node’s degree is upper bounded (approximately) by the arboricity. Additional care is taken to ensure we use the privatized arboricity in our sparsifier.
We can alternatively present our approximation guarantees in terms of $b$-matchings, which are detailed in Section 10.
Theorem 1.9 (Node-DP Approximate Maximum $b$-Matching).
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\alpha\log(n)}{\eta^{3}\varepsilon}+\frac{\log^{2}(n)}{\eta^{3}\varepsilon^{3}}+\frac{b^{\prime}\log(n)}{\eta^{2}\varepsilon}\right)$, that has the size of a $\left(2+\eta\right)$-approximate maximum $b^{\prime}$-matching.

### v3

@@ 2.1 Lower Bounds for Explicit Solutions (Section 4)
Because it is impossible to differentially privately output any subgraph (including a matching) containing only true edges, we must allow our algorithm to output “edges” that are not true edges; in other words, our matching algorithm returns a matching where only some of the edges actually exist. This corresponds to changing the input graph $G=(V,E)$ to a weighted complete graph, where every $\{u,v\}\not\in E$ is assigned weight 0 in the complete graph and actual edges have weight $1$, and asking for a max-weight matching in this graph. We can even allow our algorithm to be bicriteria and return a $b$-matching [^1] Recall that in a $b$-matching every node has degree at most $b$, thus a traditional matching is a $1$-matching. rather than a true matching. In this case, we would want our returned $b$-matching to be of comparable size to a maximum matching, while minimizing the magnitude of $b$.
We construct a family $\mathcal{M}$ of graphs over $n$ vertices with the property that for every pair of vertices $u,v$, there are neighboring graphs $G\sim G^{\prime}\in\mathcal{M}$ where $uv$ is a matching edge in $G$ but a non-edge in $G^{\prime}$. Furthermore, we show that any algorithm with reasonable utility over $\mathcal{M}$ must have non-trivial probability of outputting $uv$ (matching edge) on input $G$. Under differential privacy constraints, the probability of outputting $uv$ (non-edge) on input $G^{\prime}$ must also be non-trivial. By the symmetry of non-edge probabilities, any DP algorithm that gives a solution with decent utility must essentially output $\Omega(n^{2})$ edges. In other words, any differentially private algorithm with reasonable approximation must output many non-edges in an explicit solution.
Using generalizations of the argument above, we can in fact prove a stronger lower bound on a relaxation of the explicit matching problem to synthetic graph generation. We then show that the specific lower bound on explicit matchings stated below follows as a corollary of this more general lower bound.
@@
$e^{4\varepsilon}b+2\delta n\geq\frac{n}{2\gamma}-\beta.$
The proof of Theorem 2.1 is presented in Section 4. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then, a few meaningful regimes of $b,\gamma,\beta$ include the following:
@@
  • If $\gamma\leq\frac{\sqrt{n}}{2}$, then $\beta\geq\sqrt{n}-O(b)$. Thus, if $b\leq\sqrt{n}/c$ for some constant $c$, we still have polynomial additive loss even if we allow large polynomial multiplicative loss.
In addition to this lower bound, we also use our symmetry argument to prove the lower bounds for synthetic graph generation in Theorem 4.3 (which can be interpreted as a relaxation of explicit matchings), as well as the lower bound for implicit solutions given in Theorem 5.3.
2.2 Edge Privacy via Public Vertex Subset Mechanism (PVSM) (Sections 5 and 6)
@@
Theorem 2.1 shows that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model, there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). That is, the information released to the billboard is $\varepsilon$-differentially private and each node computes its own part of the solution by combining the public information on the billboard with its own private adjacency list. Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 3.4; such a solution has been used previously in the context of vertex cover, set cover, and maximum matching in bipartite graphs [GuptaLMRT10, hsu2014private]. It is important to note that the billboard model is not a privacy model, but a solution release model.
@@
Our results for edge privacy combine the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
@@
Although both [hsu2014private] and our results are both in the billboard model for solution release, their algorithm hinges on the “binary mechanism” [CSS11], a mechanism for releasing a stream of $n$ private counting queries with only $O(\log n)$ error, rather than the $O(\operatorname{poly}(n))$ error attained through naive composition. This mechanism releases noisy partial sums such that each query can be answered by combining $O(\log n)$ partial sums. However, this requires a centralized curator who can compute the true partial sums before adding noise, which seems incompatible with the LEDP model.
LEDP algorithm (Section 5).
Our main algorithms (Theorem 2.2 and Theorem 2.3) take inspiration from the greedy algorithms and distributed algorithms for maximal matching, although our proofs also hold for $b$-matchings. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
Within this algorithm, we introduce the Public Vertex Subset Mechanism (PVSM) which may be of independent interest to other problems such as vertex cover, dominating set, or independent set. In this mechanism, each node releases a public subset of nodes obtained via coin flips. Specifically, for each pair of nodes, we flip a set of coins with appropriate probability parameterized by subgraph indices $r$. For index $r$, each coin is flipped with probability $(1+\eta)^{-r}$. Then, a node $v$ publicly releases a subgraph index $r$ to determine a public subset of vertices consisting of coin flips which landed heads for each pair $\{v,u\}$ containing $v$. The value of $r$ determines the size of this public subset of vertices (in expectation). Then, each node $w$ takes the public subsets released by all vertices and intersects them with their private adjacency lists to determine a private subset of vertices (consisting of neighbors of $w$) only $w$ knows about. Thus, the public subset can be written on the billboard and the private subset is used to determine each node’s implicit answer. We believe this mechanism will be helpful for other graph problems that release implicit solutions. We obtain our first upper bound result for matchings in Theorem 2.2.
@@
Let $\varepsilon\in(0,1)$, $\eta\in(0,1)$, and
$b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+\Omega\left(\frac{\log(n)}{\eta^{2}\varepsilon}\right)\,.$
@@
Algorithm 2 is $\varepsilon$-LEDP and, with high probability, outputs an (implicit) $b$-matching in the billboard model that has the size of a $2$-approximate maximum $b^{\prime}$-matching.
There are two main regimes of interest, which are corollaries of Theorem 2.2:
  • If $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 is $\varepsilon$-LEDP and outputs a $b$-matching that has the size of a 2-approximate 1-matching (Corollary 5.1). Note this is a bicriteria approximation.
  • If $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$, Algorithm 2 is $\varepsilon$-LEDP and outputs a $(2+\eta)$-approximate $b$-matching (Theorem 5.2). This setting avoids the need for a bicriteria approximation.
$O(\log n)$-Round Algorithm (Section 6).
The above algorithm runs in $\Omega(n)$ rounds. To speed up our algorithm to run in polylogarithmic rounds (Theorem 2.3), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence, greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from distributed algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 2.3. Our novel utility analyses depends on a novel noisy version of a charging argument, based on edge orientation, implicitly described in [khuller1994primal]. Our improved guarantees are given in Theorem 2.3.
Theorem 2.3 (Efficient 2-Approximate Maximum $b^{\prime}$-Matching).
Let $\varepsilon\in(0,1)$, $\eta\in(0,\tfrac{1}{2})$, and matching parameter $b^{\prime}\geq 1$ where $b$ satisfies
$b\;\geq\;\frac{(1+\eta)^{2}}{1-\eta}\,b^{\prime}\;+\;\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\,\varepsilon}\right).$
@@
Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $2$-approximate maximum $b^{\prime}$-matching.
Similar to our sequential algorithm, we obtain the following results as special cases.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$, Algorithm 3 is $\varepsilon$-LEDP and outputs a $b$-matching that has the size of a 2-approximate 1-matching in $O(\log n)$ rounds (Corollary 6.1). This is a bicriteria approximation.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{{\eta^{5}}\varepsilon}\right)$, Algorithm 3 is $\varepsilon$-LEDP and outputs a $(2+\eta)$-approximate $b$-matching in $O(\log n)$ rounds (Theorem 6.2). This avoids the need for a bicriteria approximation.
Implicit Solution Lower Bound (Section 5.4).
@@
We complement our upper bounds with a lower bound which shows that even for the particular type of implicit solutions used in Theorem 2.2, our reliance on $b$ is necessary: if $b$ is sublogarithmic, then even an implicit solution cannot be a good approximation to the maximum matching in the graph. This proof also uses our symmetry argument. We start with a low-degree graph, and first use the DP guarantee over nearby graphs to argue that edges that do not exist must still be included with reasonably high probability (this is essentially a highly simplified version of the symmetry argument used in the previous lower bound). We then use group privacy to argue that there is a graph with logarithmic max degree (which is also logarithmically far from our starting graph) where many edges incident to the node of logarithmic degree must be included.
@@
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 2.2 which is the size of a $\gamma$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\gamma$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\gamma$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\gamma$, then we say that $F$ is a $\gamma$-approximate $b$-matching. with probability at least $1-\operatorname{poly}(\nicefrac{{1}}{{n}})$. Then, when decoded from the billboard, with probability at least $\nicefrac{{1}}{{2}}$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, regardless of the value of $\gamma>0$.
@@
2.3 Node Privacy via Arboricity Sparsification (Section 7)
@@
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. This model is a significantly more challenging setting than edge privacy. We use a sparsifier based on the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of our input graph for our node-DP results.
Implicit Matchings (Sections 7.1 and 7.2).
To achieve node privacy in the central model, we combine our edge privacy results with an arboricity-based sparsifier inspired by the “matching sparsifier” of [solomon2021local]. Informally, a matching sparsifier of a graph $G$ is a subgraph $H$ such that 1) the maximum matching in $H$ is approximately as large as the maximum matching in $G$, and 2) the maximum degree of $H$ is at most $O(\alpha)$, where $\alpha$ is the arboricity of $G$. Note that the average degree in a graph with arboricity $\alpha$ is at most $O(\alpha)$, so such a sparsifier essentially turns the average degree into the maximum degree while approximately preserving the maximum matching.
We give an arboricity-based sparsifier that is stable, meaning that on node-neighboring input graphs, $G$ and $G^{\prime}$, the sparsified graphs, $S$ and $S^{\prime}$, will be $\widetilde{O}(\alpha)$ apart (in edge edit distance [^4] The edge edit distance between two graphs $S$ and $S^{\prime}$ is the number of edge additions and deletions to $S$ that will create $S^{\prime}$.). Our arboricity-based sparsifier ensures the maximum degree of any vertex is $\widetilde{O}(\alpha)$. Given an algorithm to compute such a stable sparsifier, we can intuitively combine this with edge privacy to get node privacy: since the maximum degree in the sparsifier is at most $\widetilde{O}(\alpha)$ and the edge edit distance to any neighboring graph is $\widetilde{O}(\alpha)$, basic group privacy implies that we can simply use our edge-private algorithm and incur an extra loss of at most an $\widetilde{O}(\alpha)$ factor. In other words, if the maximum degree is small and the edge edit distance to any neighboring graph is also small, then node privacy is approximately the same as edge privacy, and our arboricity-based sparsifier guarantees small maximum degree bounded by the $\alpha$ of the input graph (up to polylog factors).
If we could compute and publicly release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 7.4 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 2.5).
Theorem 2.5 (Sparsifier Public Release Lower Bound).
Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-node DP algorithm which, given an input graph $G=(V,E)$, outputs an implicit matching sparsifier $E^{\prime}$. Then, if $\operatorname{OPT}(G)\leq(1+\eta)\cdot\mathbb{E}[\operatorname{OPT}(E^{\prime}\cap E)]$, the maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ even if $G$ has arboricity $1$.
Even though releasing sparsifiers does not satisfy DP, we show that we can use them in a node-DP algorithm. Intuitively, this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs. Our upper bound is given formally below.
Theorem 2.6 (Node-DP Approximate Maximum $b^{\prime}$-Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and
$b=\Omega\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}+\frac{b^{\prime}\log(n)}{\varepsilon}\right)\,.$
@@
There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model that has the size of a $\left(2+\eta\right)$-approximate maximum $b^{\prime}$-matching.
As a special case we have the following corollary:
  • If $b=\Omega\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}\right)$, there is a $\varepsilon$-node DP algorithm that outputs a $b$-matching that has the size of a $(2+\eta)$-approximate 1-matching (Corollary 7.1).
Implicit Vertex Cover (Section 7.3).
As an independent application of our arboricity sparsifier, we design an algorithm for implicit node-DP vertex cover. Similar to matchings, it is impossible to explicitly release a small vertex subset that is guaranteed to cover all edges without blatantly revealing the (non-)existence of edges. Due to this, [GuptaLMRT10] designed an $\varepsilon$-edge DP implicit $O(\nicefrac{{1}}{{\varepsilon}})$-approximate vertex cover algorithm that releases an ordering of vertices where an edge is covered by its earlier endpoint in the ordering. Arboricity sparsification also (approximately) preserves vertex cover solutions. Thus, applying our arboricity sparsifier to reduce node-DP to edge-DP with a slight modification of the [GuptaLMRT10] algorithm yields the following result.
Theorem 2.7 (Node-DP Approximate Vertex Cover).
Let $\varepsilon\in(0,1)$ and $\alpha$ denote the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit vertex cover (Definition 7.11) which is an $\widetilde{O}(\nicefrac{{\alpha}}{{\varepsilon^{2}}})$-approximation with probability $0.99$.
2.4 Utility Improvements in Bipartite Graphs over [hsu2014private] (Section 8)